OpenStudy (anonymous):

what does it mean when change in gravitational potential energy depends on: 1.mass of the object 2. Gravitational field strength 3.height which object is moved

OpenStudy (anonymous):

The equation for gravitational potential energy is: GPE = m * g * h m is the mass of the object g is the acceleration due to gravity (affected by the strength of the field) h is the height of the object Dependence just means that in order to define a quantity, those variables are necessary and appear in some form in the equation. Above, the specific relationship is that they are all multiplied together.

OpenStudy (anonymous):

ok so if mass doubles then twice the GPE right? but if it dropped on the ground you have to do work to give back GPE to the object. So why is it that if mass doubles you do twice the work?

OpenStudy (anonymous):

@Pompeii00

OpenStudy (anonymous):

Yes, if the mass doubles, there is twice the GPE. If it drops to the ground, the GPE is converted into kinetic energy (where at a height of zero, there is no remaining GPE because it is all kinetic energy). Essentially, it follows from the same equation. Work is measured in the same units as energy. It essentially IS energy from a physicist's standpoint. So if you double the mass: normal mass: GPE = mgh = GPE double mass: GPE = 2mgh = 2GPE So the work would also be double, since the energy is double. Another way to define work, more commonly, is as a Force acting over a distance. W = F * d If we consider the displacement to be the height: W = F * h Force is dependent on mass. If the mass doubles, the force must double. normal mass: W = F * h = W double mass: W = 2F * h = 2W Another way of seeing the work double due to a doubling of the mass.

OpenStudy (anonymous):

Wow hhaha Thanks!!